# FitzHugh-Nagumo equation: from global dynamics to the slow-fast system

### v1

@@ Abstract.
We concentrate our attention on the qualitative study of the phase portraits of the three-parameter FitzHugh-Nagumo family and its compactification. We divide the study into three scenarios based on the parameters. One of the scenarios is characterised by the existence of a Double-zero bifurcation with $\mathbb{Z}_{2}$–symmetry (singularity of codimension two). In this case, we explicitly exhibit the Pitchfork, Hopf, Belyakov, Double Homoclinic bifurcation/transition curves unfolding the singularity of codimension 2 and plot the bifurcation diagrams. We bridge this analysis with the theory on the associated slow-fast family and the existence of canards. We complete our study with the global compactification of the phase portraits for the family under consideration. This study complements the work summarised in Georgescu, Rocşoreanu and Giurgiţeanu, Global Bifurcations in FitzHugh-Nagumo Model, Trends in Mathematics: Bifurcations, Symmetry and Patterns (2003).
Key words and phrases:
Keywords: FitzHugh-Nagumo equation, Bifurcations, Global phase portrait, Poincaré compactification, Singular system.
[…]
@@ 1. Introduction
Information in nerve fibers is encoded through action potentials (electrical membrane changes). Alan Hodgkin and Andrew Huxley, in 1939, managed to make the first intracellular recording of an action potential by inserting microelectrodes into the giant axons of squid. In their experiments, they demonstrated that the action potentials are the result of two effects:
  • a rapid inward current carried by sodium (Na+) ions and
  • a slow activating outward current carried by potassium (K+) ions.
Hodgkin and Huxley discovered that the permeability of the membrane for $\text{Na}^{+}$ and $\text{K}^{+}$ was regulated independently, with the conductance depending on both time and the membrane potential. They provided a computational model for the action potential in a single cell, known as the Hodgkin-Huxley (HH) model [11]. Details and an overview may be found in the review [3].
@@ 1.1. From the Hodgkin-Huxley model to the FitzHugh-Nagumo equations
Hodgkin and Huxley modelled the measured changes in current by introducing precise probabilistic terms to capture ion channels, which may be either open or closed. The HH model elucidates the electrical behaviour of ion channels within the cell membrane, by addressing the passage of $\text{Na}^{+}$ and $\text{K}^{+}$ ions. The HH model consists of four ordinary differential equations, each corresponding to one of the state variables [11, Equation 1]. This system is nonlinear, making it analytically difficult to solve explicitly. Nonetheless, numerical simulations enable the exploration of certain properties and general behaviours, such as the existence of oscillations and excitability.
The fact that the HH model correctly captures these excitable dynamics was a major realisation as it provided a detailed understanding of the measured action potential and the dynamical process of neural excitability (cf. [11, Equation 1]). While the HH model effectively replicates many neuronal physiological phenomena, it s very complex. In [7], Richard FitzHugh introduced a simplified model to capture the dynamics of neuronal excitability, which was refined by Jinichi Nagumo two years later [19]. This model is known as the FitzHugh-Nagumo (FHN) model. FitzHugh focused on preserving dynamical features of the HH model, namely the presence of excitability and oscillations.
@@ 1.2. The FitzHugh-Nagumo equations
FitzHugh started from the oscillator equation (introduced in 1920) by Balthasar Van der Pol [25], that admits oscillations which are relaxation-like. This means there are periods of “low” and “high” states characterised by rapid transitions between them. Van der Pol’s equation is built from the simple differential equation for the damped harmonic oscillator. He replaced the damping constant by a damping function that depends quadratically on $x,$ thus introducing a nonlinearity ($\dot{x}$ represents the first derivative of the physical position $x$ with respect to $t$ and $c\in\mathbb{R}^{+}_{0}$):
@@
There is only “effective” damping for $|x|<1$ while for $x^{2}>1$ the nonlinear term describes amplification. To interpret the dynamics of the van der Pol equation, one can use the Liénard map [16]
@@
from which one may see the separation in time scales of both equations. For $c\gg 1$, while the first one evolves fast (order of $\mathcal{O}(c)$), the second one is much slower of the order of $\mathcal{O}(1/c),$ where $\mathcal{O}$ stands for the usual Landau notation Given a real map $f:\mathbb{R}\to\mathbb{R}$ and $c\in\mathbb{R}$, we say that $f$ is of order $\mathcal{O}(c)$ if there exist $M\in\mathbb{R}^{+}$ and $x_{0}\in\mathbb{R}$ such that $|f(x)|<Mc$, for $x>x_{0}$.. Building on the van der Pol oscillator, FitzHugh considered the equation:
@@
    the variable $x$ is related to the membrane voltage and the $\text{Na}^{+}$ activation and
@@
Equations (1.5) work as an activator-inhibitor model. Nagumo, Arimoto, and Yoshizawa proved the equivalence of this model with an electrical circuit [19]. Although the FHN model is primarily used to describe neuronal and cardiac systems, it can also be applied in a range of other biological contexts. The FHN equations provide a simplified framework for representing interconnected positive and negative feedback loops, enabling the generation of a wide range of responses such as switches, pulses, and stable oscillations [24]. Among the numerous models used to investigate cardiac cells, the FHN model is notable for being one of the simplest and most extensively researched, effectively representing the general dynamic characteristics of cardiac cells [1, 4].
@@
System (1.5) mimics an excitable system and there are limitations regarding the interpretation of the original variables of the HH model. Therefore, in the rest of the paper, we will “play safe” and stick to a mathematical motivation in terms of bifurcations and dynamical systems.
@@ 1.3. Novelty
Results in [21, pp.180] synthesised the global bifurcation diagram for the FHN model. It has been obtained by putting together as in a “huge puzzle”, all local bifurcation diagrams obtained in Chapters 2–4. See also Section 2.4 of [20]
The main novelty of this paper is the dissection of the global bifurcations diagrams of [21, pp. 180] and provide a more complete bifurcation analysis of model (1.5), namely the location of Hopf, Pitchfork and Double-zero bifurcations in some cases that will be specified in Subsection 1.4.
Besides the case-study $c=\mathcal{O}(1)$, we also study the asymptotic case $c\to\pm\infty$ ($\Leftrightarrow\varepsilon=1/c^{2}\to 0$), where canards are observed [12, 14], and we finish the analysis with the compactification of the phase portrait of (1.5) on the Poincaré disc. To the best of our knowledge, the application of this procedure to equation (1.5) is new, providing additional information about the trajectories which tend to or come from infinity. For some cases, we give the complete description of the phase portraits for (1.5) in the Poincaré disc (i.e. in the compactification of $\mathbb{R}^{2}$ adding the circle $\mathbb{S}^{1}$ of the infinity) modulo topological equivalence.
@@ 1.4. Structure
This paper is structured as follows: Section 2 introduces the definitions and preliminary concepts of blow-up and Poincaré compactification. In Section 3, we discuss the finite equilibria of (3.1) for the following three cases:
@@
Section 5 is dedicated to the study of the global phase portrait of (1.5) in the Poincaré disc. The phase portraits for each connected component in Cases A, B and C are completely illustrated. This completes the work started in [8]. Section 6 finishes the paper.
Throughout this paper, we have endeavoured to make a self-contained exposition bringing together all topics related to the proofs. We have drawn illustrative figures to make the paper easily readable and all results are illustrated with numerical simulations using the software Matlab R2015b.
2. Preparatory section
In this section, we introduce some terminology for polynomial vector fields on $\mathbb{R}^{2},$ that will be used in the remaining sections. Let $f_{(a,b,c)}$ be a smooth vector field on $\mathbb{R}^{2}$ with flow given by the unique solution $x(t)=\varphi(t,x)\in\mathbb{R}^{2}$ of the three-parameter family
$\dot{x}=f_{(a,b,c)}(x),\qquad x(0)=x_{0}\in\mathbb{R}^{2},$ (2.1)
where $\0(a,b,c)\in\mathbb{R}\times\mathbb{R}\times\mathbb{R}\backslash\{0\}$.
@@ 2.1. Useful terminology
If $Df_{(a,b,c)}$ evaluated at a given equilibrium $p$, $Df_{(a,b,c)}(p)$, has an eigenvalue with zero real part, the center manifold plays an important role and this is the set where bifurcations might occur.
For a fixed triple $\0(a,b,c)\in\mathbb{R}\times\mathbb{R}\times\mathbb{R}\backslash\{0\}$, an equilibrium point $p$ of (2.1) is called nilpotent singularity when all eigenvalues of Jacobian matrix $Df_{(a,b,c)}(p)$ are zero but $Df_{(a,b,c)}(p)\neq 0$, and is called linearly zero when $Df_{(a,b,c)}(p)\equiv 0.$ Further, the singular point $p$ is called semi-hyperbolic if exactly one eigenvalue of $Df_{(a,b,c)}(p)$ is equal to $0$
Throughout this paper, we study the Double-zero (DZ) singularity with symmetry $\mathbb{Z}_{2}$ for a family of differential equations corresponding to the case described in [27, pp. 400]. The unfolding of this singularity of codimension two involves lines of Belyakov transitions, Pitchfork, Hopf, Double Homoclinic and saddle-node of limit cycles. For the notion of super/subcritical Hopf bifurcation, we use the terminology of Golubitsky and Schaeffer [9, Chapter IV, Section 2]. We suggest the reading of [15, 18, 27] for a complete understanding of these bifurcations, as well as the sufficient conditions that prompt their existence.
@@
In this paper, we refer to an equilibrium point as undergoing a Belyakov transition if at least one pair of eigenvalues of the Jacobian matrix (evaluated at the equilibrium) changes from real to complex conjugate or vice versa, while the sign of their real part remains unchanged. Although such a transition is typically considered in higher-dimensional systems, we adopt this terminology here for the sake of brevity and clarity.
2.2. Blow-up Method
The blow-up method is one of the important methods to study the topological behaviour of solutions of a dynamical system in the neighbourhood of nilpotent and linearly zero singularities. In this method, an equilibrium point is “expanded” into a line or a circle, and new equilibria on the line or circle are examined. There are various methods for doing a blow-up. In this paper, we concentrate our attention on blow-ups in a given direction, a method which will be used in the Section 5 of the present article. In what follows, following [2], we review some concepts related to the topic.
Consider a planar polynomial differential equation associated with the vector field $\mathbf{X}=(P,Q),$ in $\mathbb{R}^{2}$ of the form:
$\left\{\begin{array}[]{l}\dot{x}_{1}=P(x_{1},x_{2})=P_{m}(x_{1},x_{2})+\cdots\\ \\ \dot{x}_{2}=Q(x_{1},x_{2})=Q_{m}(x_{1},x_{2})+\cdots\end{array}\right.$ (2.2)
where
  • $P,Q$ are coprime polynomials;
  • $P_{m},Q_{m}$ are homogeneous polynomials of degree $m\in\mathbb{N}$ and
  • the dots $\cdots$ stand for high-order terms in the variables $x_{1},x_{2}.$
Since $m>0,$ the origin is a singularity. By using the change of coordinates
$(x_{1},x_{2})\mapsto(r\cos(\theta),r\sin(\theta)),$
system (2.2) may be written in polar coordinates (if $r\neq 0$) as:
$\left\{\begin{array}[]{l}\dot{r}=\mathcal{R}(\theta)r+\cdots\\ \\ \dot{\theta}=\mathcal{T}(\theta)+\cdots\end{array}\right.$ (2.3)
where:
  • $\mathcal{R}$ and $\mathcal{T}$ are polynomials in $\cos{(\theta)}$ and $\sin{(\theta)}$ and
  • the dots $\cdots$ stand for high order terms in $r.$
If $\mathcal{T}\neq 0,$ all the solution curves tending (as $t\to\pm\infty$) to the origin are tangent to the solutions $\theta^{\star}\in[0,2\pi)$ of the equation $\mathcal{T}(\theta)=0.$ The map $\mathcal{T}$ is often called the characteristic polynomial and $\theta^{\star}$ is called characteristic direction associated with (2.2). The polynomial $\mathcal{T}$ in the cartesian coordinates $(x_{1},x_{2})$ may be written as:
$\mathcal{T}(x_{1},x_{2})=x_{1}Q_{m}(x_{1},x_{2})-x_{2}P_{m}(x_{1},x_{2}).$
@@ 2.2. Blow-up Method
Blow-up in $x_{1}$ direction is a “non-bijective” change of coordinates defined as
$(x_{1},z)\mapsto(x_{1},zx_{1})=(x_{1},x_{2}),$
where $z\geq 0$ is a new variable.
This change of coordinates projects the origin of (2.2) on the line $x_{1}=0.$ The expression of system (2.2) after the blow-up in the $x_{1}$ direction may be written as:
$\left\{\begin{array}[]{l}\dot{x}_{1}=P(x_{1},x_{1}z),\\ \\ \dot{z}=\dfrac{Q(x_{1},x_{1}z)-zP(x_{1},x_{1}z)}{x_{1}}.\end{array}\right.$ (2.4)
After the blow-up, we may cancel the common factor $x_{1}^{m-1}.$ The $x_{1}$ directional blow-up is equivalent on $x_{1}\neq 0$ and $\theta\neq\pi/2,3\pi/2,$ since there exists an analytic change of coordinates bringing $(r,\theta)$ and $(x_{1},x_{2}).$ To investigate the behavior of the solutions of system (2.2) around equilibria, it is necessary to examine the behaviour of the equilibrium points of (2.4) on the line $x_{1}=0.$ If some of these points are singular, then we may continue the blow-up process.

### v3

# slow-fast system and dynamics near infinity

@@ Abstract.
We focus on the qualitative analysis of the phase portraits arising in the three-parameter FitzHugh-Nagumo system and its compactified form. The investigation is split into three parameter-dependent cases. In one of these cases, the system displays a double-zero bifurcation with $\mathbb{Z}_{2}$–symmetry, a singularity of codimension two. For this case, we provide explicit descriptions of the bifurcation/transition curves unfolding the singularity, including Pitchfork, Hopf, Belyakov, and Double Homoclinic. Further, we present the corresponding bifurcation diagrams.
We bridge this analysis with the theory on the framework of slow-fast family and to the presence of canards.
We complete our study with the analysis of the dynamics “near infinity” for the family under consideration. This study complements the work summarised in Georgescu, Rocşoreanu and Giurgiţeanu, Global Bifurcations in FitzHugh-Nagumo Model, Trends in Mathematics: Bifurcations, Symmetry and Patterns (2003).
Alexandre A. P. Rodrigues1,† $\dagger$}Corresponding author and Nasrin Sadri2
@@ 1. Introduction
Information in nerve fibers is encoded through action potentials (electrical membrane changes). A. Hodgkin and A. Huxley, in 1939, succeeded in recording an action potential from within a neuron for the first time, using microelectrodes inserted into the squid’s giant axon. In their experiments, they demonstrated that the action potentials arise from two effects:
  • a fast inward current mediated by sodium (Na+) ions, and
  • a slow developing outward current produced by potassium (K+) ions.
Hodgkin and Huxley showed that the membrane’s permeabilities to $\text{Na}^{+}$ and $\text{K}^{+}$ are controlled separately, with their conductances varying in time and as functions of the membrane voltage. They formulated a quantitative description of the action potential in a single neuron, now referred to as the Hodgkin-Huxley (HH) model [11]. Details and an overview may be found in the review [3].
@@ 1.1. From the Hodgkin-Huxley model to the FitzHugh-Nagumo equations
Hodgkin and Huxley represented the observed current variations using probabilistic variables that describe the closing and opening of ion channels. The Hodgkin-Huxley framework characterises the electrical dynamics of these membrane channels by modelling the flow of $\text{Na}^{+}$ and $\text{K}^{+}$ ions across the membrane. In the HH framework, the dynamics are governed by four ODEs, each one associated with a specific state variable – [11, Eq. 1]. This system is nonlinear, making it analytically difficult to solve explicitly. Nonetheless, numerical simulations allow investigation of specific features and general dynamics, including the presence of excitability and oscillations.
The ability of the HH model to accurately reproduce these excitable dynamics represented a key insight, offering a complete explanation of the recorded action potential and the mechanisms underlying neural excitability (cf. [11]). While the HH model successfully reproduces many neuronal physiological phenomena, it is very complex. In [7], Richard FitzHugh introduced a simplified model to capture the dynamics of neuronal excitability, which was refined by Jinichi Nagumo two years later [19]. This system is known as the FitzHugh-Nagumo (FHN) model. FitzHugh concentrated the study on preserving the key dynamical characteristics of the HH model, specifically and oscillatory and excitability behaviour.
@@ 1.2. The FitzHugh-Nagumo equations
FitzHugh started with the oscillator equation (introduced in 1920) by Balthasar Van der Pol [25], which supports oscillations resembling relaxation-like dynamics. This means there are periods of “low” and “high” states characterised through fast transitions between them.
The Van der Pol (VdP) equation is derived from the basic differential equation describing a damped harmonic oscillator. He replaced the constant damping by a damping function that has a quadratic dependence on $x,$ resulting in the nonlinear differential equation ($\dot{x}$ denotes the first derivative of the physical position $x$ with respect to $t$ and $c\in\mathbb{R}^{+}_{0}$):
@@
The system exhibits “effective” damping only when $|x|<1$; for $x^{2}>1$, the nonlinear term leads to amplification. To analyse the dynamics of the VdP equation, the Liénard map can be employed [16]
@@
from which the time-scale separation of the equations becomes evident.
For $c\gg 1$, the first variable changes rapidly of the order of $\mathcal{O}(c)$, whereas the second evolves much more slowly, $\mathcal{O}(1/c),$ where $\mathcal{O}$ denotes the standard Landau notation (given a real map $f:\mathbb{R}\to\mathbb{R}$ and $c\in\mathbb{R}$, we say that $f$ is of order $\mathcal{O}(c)$ if there exist $M\in\mathbb{R}^{+}$ and $x_{0}\in\mathbb{R}$ such that $|f(x)|<Mc$, for $x>x_{0}$). Based on the VdP oscillator, FitzHugh considered the equation:
@@
    the variable $x$ corresponds to the membrane voltage and the $\text{Na}^{+}$ activation and
@@
Equations (1.5) act as an activator-inhibitor framework. Arimoto, Nagumo, and Yoshizawa proved that it can be realized as an electrical circuit [19]. Although the FHN model is primarily used to describe neuronal and cardiac systems, it can also be applied in a range of other biological contexts.
The FHN equations provide a simplified framework for representing interconnected negative and positive feedback loops, enabling the generation of a wide range of responses such as switches, pulses, and stable oscillations [24]. Within the wide range of models used to investigate cardiac cells, the FHN model is notable for being one of the simplest and most extensively researched, effectively representing the general dynamic characteristics of cardiac cells [1, 4].
@@
System (1.5) mimics an excitable system and there are limitations regarding the interpretation of the original variables of the HH model. In the rest of the paper, we will take a cautious approach and we stay within a mathematical motivation grounded in bifurcations and dynamical systems.
@@ 1.3. Novelty
Results in [21, pp. 180] synthesised the global bifurcation diagram for the FHN model. The result comes from integrating all the local bifurcation diagrams obtained in Chapters 2–4, as if fitting together a “huge puzzle”. See Section 2.4 of [20].
One of the novelties of this paper is the dissection of the bifurcations diagrams of [21, pp. 180] and provide a more complete bifurcation analysis of model (1.5), namely the location of Hopf, Pitchfork and Double-zero bifurcations, which will be specified in Subsection 1.4. Besides the case-study $c=\mathcal{O}(1)$, we also study the asymptotic case
$c\to\pm\infty\qquad(\Leftrightarrow\quad\varepsilon=1/c^{2}\to 0),$
where canards are observed [12, 14], and we finish the analysis with the compactification of the phase portrait of (1.5) on the Poincaré disc.
To the best of our knowledge, this procedure applied to equation (1.5) is new, providing additional information about the trajectories which come from or tend to infinity. We classify topologically all phase portraits of system (3.1) “near infinity”, represented on the Poincaré disc (i.e. within the compactification of $\mathbb{R}^{2}$ by including the circle at infinity, $\mathbb{S}^{1}$). Note that we do not claim to have derived complete phase portraits.
@@ 1.4. Structure
This paper is structured as follows: Section 2 introduces the definitions and preliminary concepts of quasi-homogeneous blow-up and Poincaré compactification. In Section 3, we discuss the finite equilibria of (3.1) for the following three cases:
@@
$a\neq 0,\,0<b<1$
Section 5 is dedicated to studying the dynamics “near infinity” for (1.5) on the Poincaré disc. The phase portraits for each connected component in Cases A, B and C are illustrated. This complements the work started in [8]. Section 6 finishes the paper.
In this paper, we aim to provide a self-contained exposition, integrating all topics pertinent to the proofs. Illustrative figures have been included for clarity, and all results are supported by numerical simulations performed using Matlab R2015b and pplane9.
2. Preparatory section
This section provides definitions for polynomial vector fields on $\mathbb{R}^{2}$ that will be used in the sections that follow. Let $f_{(a,b,c)}$ be a smooth vector field on $\mathbb{R}^{2}$ whose flow is given by the unique solution $X(t)=\varphi(t,X)\in\mathbb{R}^{2}$ of the three-parameter family
$\dot{X}=f_{(a,b,c)}(X),\qquad X(0)=X_{0}\in\mathbb{R}^{2},$ (2.1)
@@ 2.1. Useful terminology
If $Df_{(a,b,c)}$ evaluated at an equilibrium $p$, $Df_{(a,b,c)}(p)$, has an eigenvalue with zero real part, the center manifold plays an important role and this is the set where bifurcations might occur.
[…]
Throughout this paper, we study the Double-zero (DZ) singularity with symmetry $\mathbb{Z}_{2}$ for a family of differential equations corresponding to the case described in [27, pp. 400]. The unfolding of this singularity of codimension two involves lines of Belyakov transitions, Pitchfork, Hopf, Double Homoclinic and saddle-node bifurcation of limit cycles.
A Hopf bifurcation is termed supercritical when the first Lyapunov coefficient of its normal form is negative, leading to the emergence of a stable small-amplitude limit cycle, and subcritical when this coefficient is positive, in which an unstable small-amplitude limit cycle is created [26]. For full details on these bifurcations and the criteria that guarantee their appearance, see [15, 18, 27].
@@
In this paper, we refer to an equilibrium point $p$ as undergoing a Belyakov transition if at least one pair of eigenvalues of $Df_{(a,b,c)}(p)$ changes from real to complex conjugate or vice versa, while the sign of their real part remains unchanged. Although such a transition is typically considered in higher-dimensional systems, we adopt this terminology here for the sake of brevity and clarity.
2.2. Quasi-homogeneous blow-up method
Quasi-homogeneous blow-ups are used to analyze and simplify the behaviour of differential equations near non-hyperbolic singularities by rescaling the variables according to their natural weights. This technique transforms the singularity into a more regular, analyzable object. Although homogeneous blow-ups
$\phi:\mathbb{S}^{1}\times\mathbb{R}\to\mathbb{R}^{2},\quad(\theta,\rho)\mapsto\left(\rho\cos\theta,\,\rho\sin\theta\right),$
suffice for the study of isolated singularities of an analytic vector field, the quasi-homogeneous approach (scaling the variables with different degrees) is considerably more efficient. Here we review this method briefly – for more details and examples, see [5, Section 3.3].
Consider the polynomial vector field $f=(P,Q)$ associated with the planar differential equation:
$\left\{\begin{array}[]{l}\dot{x}_{1}=P(x_{1},x_{2})\\ \\ \dot{x}_{2}=Q(x_{1},x_{2}).\end{array}\right$. (2.2)
In the language of singularity theory, we have:
$f(x_{1},x_{2})=P(x_{1},x_{2})\frac{\partial}{\partial x_{1}}+Q(x_{1},x_{2})\frac{\partial}{\partial x_{2}},$
with an isolated singular point at the origin $(0,0)$. Since $P$ and $Q$ are polynomial we may find $p_{ij},q_{ij}\in\mathbb{R}$ such that:
$P(x_{1},x_{2})=\sum_{i+j\geq 1}p_{ij}x_{1}^{i}x_{2}^{j}\qquad\text{and}\qquad Q(x_{1},x_{2})=\sum_{i+j\geq 1}q_{ij}x_{1}^{i}x_{2}^{j}.$
@@ 2.2. Quasi-homogeneous blow-up method
The quasi-homogeneous blow-up is defined by
$\phi:\mathbb{S}^{1}\times\mathbb{R}\to\mathbb{R}^{2},\quad(\theta,\rho)\mapsto\left(\rho^{\alpha}\cos\theta,\,\rho^{\beta}\sin\theta\right)=(x_{1},x_{2}),$
for some $\alpha,\beta\in\mathbb{N}.$
If $(\alpha,\beta)=(1,1)$, this corresponds to the standard homogeneous case. Doing the quasi-homogeneous blow-up (of the previous definition) on $f=(P,Q)$, it yields the following differential equation:
$\dot{\rho}=\frac{P\rho^{\beta}\cos\theta+Q\rho^{\alpha}\sin\theta}{\rho^{\alpha+\beta-1}(\alpha\cos^{2}\theta+\beta\sin^{2}\theta)},\qquad\dot{\theta}=\frac{\alpha Q\rho^{\alpha}\cos\theta-\beta P\rho^{\beta}\sin\theta}{\rho^{\alpha+\beta}(\alpha\cos^{2}\theta+\beta\sin^{2}\theta)},$
in which $P$ and $Q$ are computed at $(\rho^{\alpha}\cos\theta,\rho^{\beta}\sin\theta)$. A common factor $\rho^{d}>0$ can be cancelled for a suitable choice of $d>0$.
Throughout this article, we will be interested in quasi-homogeneous directional blow-ups in a given direction. This technique isolates and desingularizes the dynamics of trajectories approaching the singularity along that direction, creating a chart where the vector field becomes simpler and its behaviour can be analyzed separately from other directions.
Definition 2.3.
The quasi-homogeneous directional blow-up along the positive (resp. negative) $x$ axis is given by the map:
$\overline{f}^{x}_{+}:\quad(\bar{x},\bar{y})\mapsto(\bar{x}^{\alpha},\bar{x}^{\beta}\bar{y}),$
$\overline{f}^{x}_{-}:\quad(\bar{x},\bar{y})\mapsto(-\bar{x}^{\alpha},\bar{x}^{\beta}\bar{y}),$
and the quasi-homogeneous directional blow-up along the positive (resp. negative) $y$ axis is described by:
$\overline{f}^{y}_{+}:\quad(\bar{x},\bar{y})\mapsto(\bar{x}\bar{y}^{\alpha},\bar{y}^{\beta}),$
$\overline{f}^{y}_{-}:\quad(\bar{x},\bar{y})\mapsto(\bar{x}\bar{y}^{\alpha},-\bar{y}^{\beta}),$
where $\bar{x}$ and $\bar{y}$ denote new variables and $\alpha,\beta\in\mathbb{N}.$
Among all possible pairs $(\alpha,\beta)\in\mathbb{N}\times\mathbb{N}$ of the definition of quasi-homogeneous blow-ups, a natural question is which choice is the most “efficient” (in the sense that we desingularize the equilibria in the minimum number of steps) for performing a quasi-homogeneous blow-up to simplify and resolve a singularity. Selecting such an efficient pair balances the relative importance of the variables, making the leading-order dynamics near the singularity clear. The Newton polygon provides a systematic tool to guide this choice, which we now introduce (adapted from [5]).
Definition 2.4.
We define the support of $f$ as the subset
$S=\{(m-1,n)\mid p_{mn}\neq 0\}\cup\{(m,n-1)\mid q_{mn}\neq 0\}\subset\mathbb{R}^{2},$
where the point (1,n)(-1,n) corresponds to the monomial $p_{0n}x_{2}^{n}$ and the point (m,1)(m,-1) corresponds to the monomial $q_{m0}x_{1}^{m}.$ Furthermore, $(0,0)$ is associated with the monomials $p_{10}x_{1}$ and $q_{01}x_{2}.$
Definition 2.5.
The Newton polygon associated to $f$ is given by the convex hull
$\bigcup_{(r,s)\in S}\{({r}^{*},{s}^{*})\in\mathbb{R}^{2}\mid r\leq{r}^{*},\,s\leq{s}^{*}\}.$
The Newton polygon associated to $f$ consists of the union $\gamma$ of the points/segments $\gamma_{k}$, enumerated in order from left to right. For each $1\leq k\leq n$, the segment $\gamma_{k}$ lies on the straight line $\alpha_{k}r+\beta_{k}s=d_{k}$ where $\alpha_{k}$ and $\beta_{k}$ are coprime integers and $d_{k}$ is a constant. Since $f$ has an isolated singularity at the origin, then at least one of the two points $(0,s)$ or (1,s)(-1,s) belongs to the support $S$ for some $s$ and, similarly, at least one of the two points (r,1)(r,-1) or $(r,0)$ lies in $S$, for some $r$. Thus a first segment $\gamma_{1}$ always exists in the Newton diagram ($\Leftrightarrow$ it is non-empty).
The Newton diagram identifies the dominant monomials near a singularity; the slopes of its compact edges determine the weight ratio $\beta/\alpha$, $\alpha\neq 0$, used to choose the exponents of a quasi-homogeneous blow-up.
Proposition 2.6.
[5, Section 3.3, adapted]. Under the previous conditions, the most efficient pair $(\alpha,\beta)$ for the quasi-homogeneous blow-up corresponds to the edge of the Newton polygon with minimal slope (in absolute value).